complete the square, and find the roots of the quadratic equation.
x^2 + 16x = 0

Answer by Prithwis(166)

x^2 + 16x = 0
=> x^2 + 16x + 64 = 64
=> (x+8)^2 = 8^2
=> (x+8) = 8 or (x+8) = -8
=> x = 0 or x = -16